# Answers to frequently asked questions about the pulsar timing array Hellings and Downs curve

### v2

@@ 2.4 Visualizing the redshift response (antenna pattern functions)
To end this section, we leave it as an exercise for the reader to calculate the values of the various antenna pattern functions $|G(\hat{\Omega})|$, $|F(\hat{\Omega})|$, and ${\cal R}(f,\hat{\Omega})$ for the different pulsar locations and GW shown in Figure 7. Similar to Figure 2, there is a single GW point source in direction $-\hat{\Omega}$ relative to Earth, sufficiently far away so that the waves are planar when they reach the vicinity of the Earth and the pulsars. We also assume that the GW is monochromatic with $fL\equiv L/\lambda=5$, where $\lambda$ is the GW wavelength and $L$ is the distance of the pulsars from Earth. The values of the various antenna patterns are given in Table 1.
@@ 2.5 The plane wave expansion
[…]
Here and elsewhere, frequency and time integrals are over the range (,)($\infty$,$\infty$)( - ∞ , ∞ ), unless explicit limits are given.
@@
Substitute this plane wave expansion (1cdmal) into the formula for the redshift (1cdmn), differentiate with respect to $t$, and then integrate along the null ray coming from pulsar “a” as specified in (1cdmb). One obtains a compact formula for the redshift $Z_{a}(t)$ of an arbitrary pulsar in an arbitrary GW background:
[…]
@@
$R_{a}^{A}(f,{\hat{\Omega}})\equiv\left[1-e^{-i2\pi fL_{a}(1+\hat{\Omega}\cdot\hat{p}_{a})}\right]F_{a}^{A}(\hat{\Omega})\,\ \Leftrightarrow\ R_{a}=R_{a}^{+}+iR_{a}^{\times}$ (1cdman)
@@
$F_{a}^{A}(\hat{\Omega})\equiv\frac{1}{2}\frac{{\hat{p}_{a}}^{i}{\hat{p}_{a}}^{j}}{1+{\hat{\Omega}}\cdot{\hat{p}_{a}}}e_{ij}^{A}(\hat{\Omega})\,,\quad A=+,\times\,.$ (1cdmao)
@@ 3 Pulsar-pair correlations
Large amplitude GWs would be directly visible in the GW detector output. Indeed, in the early days of pulsar timing, upper limits were set on the amplitude of GWs by examining the timing residuals from single pulsars. The smallness of those residuals meant that the GW amplitudes could not have been too large.
However, the GWs predicted for typical background sources are too weak to be directly visible in the data; they cannot be separated from noise in the measurements or other physical effects. Fortunately, there is a way around this: look for a common GW signal by correlating the redshift measurements for two or more Earth-pulsar baselines for a PTA. With sufficient integration or averaging, one can “dig below the noise”. The same method (correlating signals from pairs of LIGO-like Earth-based detectors) is used to hunt for a stochastic background signal at audio frequencies [16, 10, 6].
@@ 3.1 Hellings and Downs correlation (source averaging)
$\rho_{ab}\equiv\overline{Z_{a}(t)Z_{b}(t)}\equiv\frac{1}{T}\int_{-T/2}^{T/2}{\rm d}t\>Z_{a}(t)Z_{b}(t)\,.$ (1cdmap)
@@
The frequency-domain representation for a single GW source is a special case of the general plane wave expansion (1cdmal), where the Fourier amplitudes vanish except for a single direction $\hat{\Omega}$. Alternatively, and equivalently, we can return to the formulation (1a) for a single source, and define Fourier amplitudes $\tilde{h}^{A}(f)$ for the source waveforms by
$\tilde{h}^{A}(f)\equiv\int{\rm d}t\>e^{-i2\pi ft}\,h^{A}(t)\ \Leftrightarrow\ \tilde{h}(f)=\tilde{h}^{+}(f)+i\tilde{h}^{\times}(f)\,,$ (1cdmaq)
@@
The right-hand side of (1cdmap) contains 16 terms, because each pulsar redshift has contributions from two polarization components evaluated at both the Earth and pulsar, so there are four terms per pulsar. All of these terms contribute to $\rho_{ab}$ for a fixed set of GW sources and a fixed pair of pulsars.
However, for calculating the mean correlation for a pair of distinct Earth-pulsar baselines, only four terms survive, as described in detail in, e.g., [8, 9]. The argument that one uses to show this depends on the averaging procedure used to define the mean. The first form of averaging, employed by Hellings and Downs [5], keeps the pulsars fixed and averages the correlation $\rho_{ab}$ over GW source directions $\hat{\Omega}$, assuming that the sources are unpolarized and isotropic on the sky.
By performing the all-sky average for an unpolarized signal, the Earth-pulsar and pulsar-pulsar terms enter the expected correlation via integrals over $\hat{\Omega}$ of terms proportional to the auto-correlation function $C(\tau)$ of the GW signal [^5] The auto-correlation function $C(\tau)$ is the Fourier transform of the power spectrum of the GW signal, $C(\tau)\equiv\int{\rm d}f\>e^{i2\pi f\tau}|\tilde{h}(f)|^{2}/T$. evaluated at $\tau=\tau_{a}$, $\tau_{b}$, and $\tau_{a}-\tau_{b}$, where $\tau_{a}\equiv L_{a}(1+\hat{\Omega}\cdot\hat{p}_{a})$ and $\tau_{b}\equiv L_{b}(1+\hat{\Omega}\cdot\hat{p}_{b})$. As shown in [9], if we assume that the coherence length of the GW signal is short compared to $\tau_{a}$, $\tau_{b}$, and $|\tau_{a}-\tau_{b}|$, then these terms can be ignored relative to the Earth-Earth terms [see (C13) and Figures 8 and 9 in Appendix C of [9] for details]. However, for the correlation of a single pulsar with itself (i.e., $a=b$), it immediately follows that $\tau_{a}-\tau_{b}=0$. Thus, the auto-correlation function is evaluated at 0, and hence doubles the contribution of the Earth-Earth term for a stationary GW background.
@@
$\langle\rho_{ab}\rangle=h^{2}\mu_{\rm u}(\gamma)[1+\delta_{ab}]\,,$ (1cdmas)
@@
$h^{2}\equiv\frac{1}{2}\left(\overline{(h^{+})^{2}+(h^{\times})^{2}}\right)=\frac{1}{2}\overline{|h|^{2}}\,,$ (1cdmata)
$\mu_{\rm u}(\gamma)\equiv\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>\sum_{A}F^{A}_{a}(\hat{\Omega})F^{A}_{b}(\hat{\Omega})=\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>F_{a}(\hat{\Omega})F^{*}_{b}(\hat{\Omega})\,.$ (1cdmatb)
Here, $h^{2}$ is the squared GW strain and $\mu_{\rm u}(\gamma)$ is Hellings and Downs curve, which is the mean Hellings and Downs correlation for an unpolarized and isotropic GW background [5]. There is no contribution from the cross-polarization terms since $\overline{h^{+}h^{\times}}=0$ for unpolarized GW sources. The $\delta_{ab}$ term in (1cdmas) is needed to account for auto-correlations.
Simple symmetry arguments show that the integral in (1cdmatb) can depend only on the angle $\gamma$ between the lines of sight to the two pulsars. So, without loss of generality, we can place pulsar $a$ on the $z$-axis and pulsar $b$ in the $xz$-plane so that $\hat{p}_{a}=\hat{z}$ and $\hat{p}_{b}=\sin\gamma\,\hat{x}+\cos\gamma\,\hat{z}$. The polarization tensors needed to calculate $F^{A}_{a}(\hat{\Omega})$ and $F^{A}_{b}(\hat{\Omega})$ are defined by (1ca), (1cb), with $\hat{\Omega}$, $\hat{m}$, $\hat{n}$ given by (1cda), (1cdb), (1cdc). The integral can be evaluated analytically using contour integration, as detailed in [8, 9, 17], giving
$\mu_{\rm u}(\gamma)=\frac{1}{3}-\frac{1}{6}\left(\frac{1-\cos\gamma}{2}\right)+\left(\frac{1-\cos\gamma}{2}\right)\ln\left(\frac{1-\cos\gamma}{2}\right)\,.$ (1cdmatau)
@@
We leave it as an exercise for the reader to show that a similar calculation of the expected correlation for an array of “short-arm” LIGO-like detectors yields [^6] We have used a proportionality sign in (1cdmatav) since there are frequency-dependent factors of $2\pi fL$ in (1cdmah) that we are ignoring.
$\mu_{\rm u}(\gamma)\propto P_{2}(\cos\gamma)\,,$ (1cdmatav)
where $P_{2}(x)$ is the $l=2$ Legendre polynomial. For this calculation, one needs to replace the response functions $R^{A}(f,\hat{\Omega})$ in (1cdmaf) with the short-arm limiting expressions (1cdmah). This substitution simplifies the integrand in (1cdmatb) to quadratics in $\sin\phi$ and $\cos\phi$ and quartics in $\cos\theta$. The short-arm expected correlation is plotted in the right panel of Figure 9. Since the short-arm function involves only the $l=2$ Legendre polynomial, it is an example of a purely quadrupolar correlation.
Figure 9: Left panel: Expected Hellings and Downs correlation for pairs of PTA pulsars. Right panel: Expected correlation for pairs of short-arm LIGO-like detectors, normalized with the same zero-angular-separation value of 1/3. Note that the zero-angular-separation value in the left panel corresponds to a pair of pulsars located close on the sky but at significantly different distances. If they are much closer than a GW wavelength, or if they are the same pulsar, then the correlation equals twice the zero-angular-separation value that is shown in this plot.
@@ 3.2 Pulsar averaging
The second form of averaging, originally proposed by Cornish and Sesana [18] and then further developed by the authors in [9, 19], is called pulsar averaging. Here, for a fixed set of GW sources, one averages the correlation $\rho_{ab}$ over all pulsar pairs separated by the same angle $\gamma$, assuming that one has access to a large number of pulsar pairs uniformly distributed on the sky. Pulsar averaging, like source averaging, eliminates the pulsar-pulsar and Earth-pulsar terms, provided the distances to pulsars in the same direction $\hat{p}$ are large compared to the GW wavelengths that PTAs are expected to detect. The result of pulsar averaging is a quantity
$\Gamma(\gamma)\equiv\langle\rho_{ab}\rangle_{ab\in\gamma}\,,$ (1cdmataw)
@@
Pulsar averaging corresponds to observational practice. PTA collaborations monitor many pulsars distributed across the sky, so they can average together the correlations from all pulsar pairs lying in an angular separation bin centered on angle $\gamma$. As more pulsars are added to a PTA, the closer this “binned” pulsar averaging will get to ideal pulsar averaging, where one has access in principle to an infinite number of pulsar pairs uniformly distributed over the sky. In contrast, source averaging is not observationally possible, since we have access to only one Universe and its associated (fixed) collection of GW sources.
As shown in [9, 18], pulsar averaging yields the mean correlation (1cdmas) as the original Hellings and Downs source-averaging prescription for a single GW source:
$\Gamma(\gamma)=h^{2}\mu_{\rm u}(\gamma)[1+\delta_{ab}]\,.$ (1cdmatax)

### v3

@@ 2.4 Visualizing the redshift response (antenna pattern functions)
These differences in antenna patterns and responses for “short-arm” and “long-arm” GW detectors have close equivalents in the context of electromagnetic (EM) antennas and phased arrays. Those which are small compared to the EM wavelength are analogous to short-arm GW detectors, and vice versa. Because of reciprocity, it does not matter if the antennas are receiving or broadcasting.
A simple example is the center-fed linear antenna. This a straight wire of length $d$, laid along the $z$-axis between $-d/2<z<d/2$, and broken open at the origin, where the two halves are driven by a voltage difference that varies sinusoidally with time. This creates a current distribution along the wire which is sinsoidal with respect to both time and space, and which vanishes at the ends of the wire, see [18, Sec. 9.4] for details. In the radiation zone $r\gg d$, $r\gg\lambda$, where $r$ is the distance from the antenna, the electric field $\vec{E}$ has an azimuthally-symmetric magnitude proportional to
$|\vec{E}|\propto\frac{\cos\left(\frac{\pi d}{\lambda}\cos\theta\right)-\cos\left(\frac{\pi d}{\lambda}\right)}{\sin\theta}\frac{1}{r}\,.$ (1cdmal)
If the antenna is long, $d\gg\lambda$, then this radiation pattern has multiple lobes, similar to that shown in the rightmost panel of Figure 6 for a PTA. If the antenna is short, $d\ll\lambda$, then (1cdmal) approaches the single-lobed radiation pattern of a short EM dipole antenna $r|\vec{E}|\propto(\pi d/\lambda)^{2}\,\sin\theta$. This is similar to the response shown in the left-hand panel of Figure 6 for a short-arm GW detector.
In the EM case, the multiple lobes are due to constructive and destructive interference from the sinusoidally-varying current distribution along the length of the long antenna. In the GW PTA case, the multiple lobes are due to constructive and destructive interference of the sinusoidally-varying GW amplitude along the photon path.
To end this section, we leave it as an exercise for the reader to calculate the values of the various antenna pattern functions $|G(\hat{\Omega})|$, $|F(\hat{\Omega})|$, and ${\cal R}(f,\hat{\Omega})$ for the different pulsar locations and GW shown in Figure 7. As in Figure 2, there is a single GW point source in direction $-\hat{\Omega}$ relative to Earth, sufficiently far away that the wavefronts are planar when they reach the vicinity of the Earth and the pulsars. We also assume that the GW is monochromatic with $fL\equiv L/\lambda=5$, where $\lambda$ is the GW wavelength and $L$ is the distance of the pulsars from Earth. The values of the various antenna patterns are given in Table 1.
@@ 2.5 The plane wave expansion
$h_{ij}(t,\vec{x})=\sum_{A}\int{\rm d}f\int{\rm d}\hat{\Omega}\,h_{A}(f,\hat{\Omega})e^{A}_{ij}(\hat{\Omega}){\rm e}^{i2\pi f(t-\hat{\Omega}\cdot\vec{x})}\,,$ (1cdmam)
where $\int{\rm d}\hat{\Omega}\equiv\int_{0}^{2\pi}{\rm d}\phi\int_{0}^{\pi}\sin\theta\,{\rm d}\theta$. Here and elsewhere, frequency and time integrals are over the range (,)($\infty$,$\infty$)( - ∞ , ∞ ), unless explicit limits are given.
@@
Substitute this plane wave expansion (1cdmam) into the formula for the redshift (1cdmn), differentiate with respect to $t$, and then integrate along the null ray coming from pulsar “a” as specified in (1cdmb). One obtains a compact formula for the redshift $Z_{a}(t)$ of an arbitrary pulsar in an arbitrary GW background:
$Z_{a}(t)$ $=$ $\sum_{A}\int{\rm d}f\int{\rm d}\hat{\Omega}\,R_{a}^{A}(f,{\hat{\Omega}})h_{A}(f,\hat{\Omega}){\rm e}^{i2\pi ft}\,.$ (1cdman)
@@
$R_{a}^{A}(f,{\hat{\Omega}})\equiv\left[1-e^{-i2\pi fL_{a}(1+\hat{\Omega}\cdot\hat{p}_{a})}\right]F_{a}^{A}(\hat{\Omega})\,\ \Leftrightarrow\ R_{a}=R_{a}^{+}+iR_{a}^{\times}$ (1cdmao)
@@
$F_{a}^{A}(\hat{\Omega})\equiv\frac{1}{2}\frac{{\hat{p}_{a}}^{i}{\hat{p}_{a}}^{j}}{1+{\hat{\Omega}}\cdot{\hat{p}_{a}}}e_{ij}^{A}(\hat{\Omega})\,,\quad A=+,\times\,.$ (1cdmap)
@@ 3 Pulsar-pair correlations
Large amplitude GWs would be directly visible in the GW detector output. Indeed, in the early days of pulsar timing, upper limits were set on the amplitude of GWs by examining the timing residuals from single pulsars [19]. The smallness of those residuals meant that the GW amplitudes could not have been too large.
However, the GWs predicted for typical background sources are too weak to be directly visible in the data; they cannot be separated from noise in the measurements or other physical effects. Fortunately, there is a way around this: look for a common GW signal by correlating the redshift measurements for two or more Earth-pulsar baselines for a PTA. With sufficient integration or averaging, one can “dig below the noise”. The same method (correlating signals from pairs of LIGO-like Earth-based detectors) is used to hunt for a stochastic GW background at audio frequencies [20, 10, 7].
@@ 3.1 Hellings and Downs correlation (source averaging)
$\rho_{ab}\equiv\overline{Z_{a}(t)Z_{b}(t)}\equiv\frac{1}{T}\int_{-T/2}^{T/2}{\rm d}t\>Z_{a}(t)Z_{b}(t)\,.$ (1cdmaq)
@@
The frequency-domain representation for a single GW source is a special case of the general plane wave expansion (1cdmam), where the Fourier amplitudes vanish except for a single direction $\hat{\Omega}$. Alternatively, and equivalently, we can return to the formulation (1a) for a single source, and define Fourier amplitudes $\tilde{h}^{A}(f)$ for the source waveforms by
$\tilde{h}^{A}(f)\equiv\int{\rm d}t\>e^{-i2\pi ft}\,h^{A}(t)\ \Leftrightarrow\ \tilde{h}(f)=\tilde{h}^{+}(f)+i\tilde{h}^{\times}(f)\,,$ (1cdmar)
@@
The right-hand side of (1cdmaq) contains 16 terms, because each pulsar redshift has contributions from two polarization components evaluated at both the Earth and pulsar, so there are four terms per pulsar. All of these terms contribute to $\rho_{ab}$ for a fixed set of GW sources and a fixed pair of pulsars.
However, for calculating the mean correlation for a pair of distinct Earth-pulsar baselines, only four terms survive, as described in detail in, e.g., [9, 5]. The argument that one uses to show this depends on the averaging procedure used to define the mean.
The first form of averaging, employed by Hellings and Downs [6], keeps the pulsars fixed and averages the correlation $\rho_{ab}$ over GW source directions $\hat{\Omega}$, assuming that the sources are unpolarized and isotropic on the sky. In doing this, Hellings and Downs also (implicitly) assumed that distinct sources were uncorrelated, because they dropped cross-terms arising from the interference between different GW sources. This is valid if one computes an average over many different universes, but such terms are present in any given universe and cause deviations from that average, as discussed in Question 7.
By performing the all-sky average for an unpolarized signal, the Earth-pulsar and pulsar-pulsar terms enter the expected correlation via integrals over $\hat{\Omega}$ of terms proportional to the auto-correlation function $C(\tau)$ of the GW signal [^7] The auto-correlation function $C(\tau)$ is the Fourier transform of the power spectrum of the GW signal, $C(\tau)\equiv\int{\rm d}f\,e^{i2\pi f\tau}|\tilde{h}(f)|^{2}\big{/}\int{\rm d}f^{\prime}\,|\tilde{h}(f^{\prime})|^{2}$. evaluated at $\tau=\tau_{a}$, $\tau_{b}$, and $\tau_{a}-\tau_{b}$, where $\tau_{a}\equiv L_{a}(1+\hat{\Omega}\cdot\hat{p}_{a})$ and $\tau_{b}\equiv L_{b}(1+\hat{\Omega}\cdot\hat{p}_{b})$. As shown in [5], if we assume that the coherence time of the GW signal is short compared to $\tau_{a}$, $\tau_{b}$, and $|\tau_{a}-\tau_{b}|$, then these terms can be ignored relative to the Earth-Earth terms [see (C13) and Figures 8 and 9 in Appendix C of [5] for details]. However, for the correlation of a single pulsar with itself (i.e., $a=b$), it immediately follows that $\tau_{a}-\tau_{b}=0$. Thus, the auto-correlation function is evaluated at 0, and hence doubles the contribution of the Earth-Earth term for a stationary GW background.
@@
$\langle\rho_{ab}\rangle=h^{2}\mu_{\rm u}(\gamma)[1+\delta_{ab}]\,,$ (1cdmat)
@@
$h^{2}\equiv\frac{1}{2}\left(\overline{(h^{+})^{2}+(h^{\times})^{2}}\right)=\frac{1}{2}\overline{|h|^{2}}\,,$ (1cdmaua)
$\mu_{\rm u}(\gamma)\equiv\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>\sum_{A}F^{A}_{a}(\hat{\Omega})F^{A}_{b}(\hat{\Omega})=\frac{1}{4\pi}\int{\rm d}\hat{\Omega}\>F_{a}(\hat{\Omega})F^{*}_{b}(\hat{\Omega})\,.$ (1cdmaub)
Here, $h^{2}$ is the squared GW strain and $\mu_{\rm u}(\gamma)$ is the Hellings and Downs curve, which is the mean Hellings and Downs correlation for an unpolarized and isotropic GW background [6]. There is no contribution from the cross-polarization terms since $\overline{h^{+}h^{\times}}=0$ for unpolarized GW sources. The $\delta_{ab}$ term in (1cdmat) is needed to account for auto-correlations.
Simple symmetry arguments show that the integral in (1cdmaub) can depend only on the angle $\gamma$ between the lines of sight to the two pulsars. So, without loss of generality, we can place pulsar $a$ on the $z$-axis and pulsar $b$ in the $xz$-plane so that $\hat{p}_{a}=\hat{z}$ and $\hat{p}_{b}=\sin\gamma\,\hat{x}+\cos\gamma\,\hat{z}$. The polarization tensors needed to calculate $F^{A}_{a}(\hat{\Omega})$ and $F^{A}_{b}(\hat{\Omega})$ are defined by (1ca), (1cb), with $\hat{\Omega}$, $\hat{m}$, $\hat{n}$ given by (1cda), (1cdb), (1cdc). The integral can be evaluated analytically using contour integration, as detailed in [9, 5, 21], giving
$\mu_{\rm u}(\gamma)=\frac{1}{3}-\frac{1}{6}\left(\frac{1-\cos\gamma}{2}\right)+\left(\frac{1-\cos\gamma}{2}\right)\ln\left(\frac{1-\cos\gamma}{2}\right)\,.$ (1cdmauav)
@@
We leave it as an exercise for the reader to show that a similar calculation of the expected correlation for an array of “short-arm” LIGO-like detectors yields [^8] We have used a proportionality sign in (1cdmauaw) since there are frequency-dependent factors of $2\pi fL$ in (1cdmah) that we are ignoring.
$\mu_{\rm u}(\gamma)\propto P_{2}(\cos\gamma)=\frac{1}{2}(3\cos^{2}\gamma-1)\,,$ (1cdmauaw)
where $P_{2}(x)$ is the $l=2$ Legendre polynomial. For this calculation, one needs to replace the response functions $R^{A}(f,\hat{\Omega})$ in (1cdmaf) with the short-arm limiting expressions (1cdmah). This substitution simplifies the integrand in (1cdmaub) to quadratics in $\sin\phi$ and $\cos\phi$ and quartics in $\cos\theta$. The short-arm expected correlation is plotted in the right panel of Figure 9. Since the short-arm function involves only the $l=2$ Legendre polynomial, it is an example of a purely quadrupolar correlation.
Figure 9: Left panel: Expected Hellings and Downs correlation for pairs of PTA pulsars. Right panel: Expected correlation for pairs of short-arm LIGO-like detectors, normalized with the same zero-angular-separation value of 1/3. Note that the zero-angular-separation value in the left panel corresponds to a pair of pulsars located close on the sky but at significantly different distances. If they are much closer than a GW wavelength, or if they are the same pulsar, then the correlation is twice the zero-angular-separation value that is shown in this plot.
@@ 3.2 Pulsar averaging
The second form of averaging, originally employed by Cornish and Sesana [22] and then further developed by the authors in [5, 23], is called pulsar averaging. Here, for a fixed set of GW sources, one averages the correlation $\rho_{ab}$ over all pulsar pairs separated by the same angle $\gamma$, assuming that one has access to a large number of pulsar pairs uniformly distributed on the sky. Pulsar averaging, like source averaging, eliminates the pulsar-pulsar and Earth-pulsar terms, provided the distances to pulsars in the same direction $\hat{p}$ are large compared to the GW wavelengths that PTAs are expected to detect. The result of pulsar averaging is a quantity
$\Gamma(\gamma)\equiv\langle\rho_{ab}\rangle_{ab\in\gamma}\,,$ (1cdmauax)
@@
Pulsar averaging corresponds to observational practice. PTA collaborations monitor many pulsars distributed across the sky, so they can average together the correlations from all pulsar pairs lying in an angular separation bin centered on angle $\gamma$. As more pulsars are added to a PTA, this “binned” pulsar averaging will get closer to ideal pulsar averaging, where one has access in principle to an infinite number of pulsar pairs uniformly distributed over the sky. In contrast, source averaging is not observationally possible, since we have access to only one Universe and its associated (fixed) collection of GW sources.
As shown in [5, 22], pulsar averaging yields the same mean correlation (1cdmat) as the original Hellings and Downs source-averaging prescription for a single GW source:
$\Gamma(\gamma)=h^{2}\mu_{\rm u}(\gamma)[1+\delta_{ab}]\,.$ (1cdmauay)